OpenStudy (anonymous):

On the function y=5x^2-2, what effect does the number 5 have on the graph, as compared to the graph of y=x^2? A. It stretches the graph vertically by a factor of 5. B. It stretches the graph horizontally by a factor of 5. C. It shrinks the graph horizontally by a factor of 5. D. It shrinks the graph vertically by a factor of 5.

OpenStudy (jdoe0001):

\(\bf y = A ( Bx + C ) + D\) A = stretch or expand factor D = vertical shift, D > 0, up, D < 0 down

OpenStudy (anonymous):

ummmm

OpenStudy (jdoe0001):

\(\large \begin{array}{cllll} y = {\color{blue}{ A}} ( Bx + C ) + {\color{brown}{ D}} \\ \quad \\ y={\color{blue}{ 5}}x^2{\color{brown}{ -2}} \end{array}\)

OpenStudy (anonymous):

ok